OpenStudy (angelax0):

Write the quadratic equation whose roots are −4 and 6, and whose leading coefficient is 5.

OpenStudy (mathmale):

You might begin with the general form of the quadratic equation, which is \[ax^2+bx+c=0\]

OpenStudy (mathmale):

In terms of this equation, what can you say if you're told that -4 and 6 are roots of the equation and that the leading coefficient is 5? Write this equation out, once for the root -4 and once for the root 6. These two equations provide you with enuf info to determine the values of the coefficients b and c.

OpenStudy (angelax0):

so would it be 5x^2+20x-30=0?

OpenStudy (mathmale):

You need to replace x by either -4 or 6. This will eliminate x from the equation. Yor goal is to find b and c. If ax^2+bx+c=0, and x=-4, then 5(-4)^2+b(-4) +c=0. Write a similar equation for the root x=6. show your work. simplify your results. Solve your two resulting equations for b and c.

OpenStudy (angelax0):

write*

OpenStudy (mathmale):

Actually, you ARE trying to solve for b and c. Please follow my suggestions.

OpenStudy (mathmale):

Coefficient a has the value 5 (given) Find coefficients b and c. Then write out your quadratic: y = ax^2 + bx + c.

OpenStudy (zzr0ck3r):

\(y=a(x-b)(x-c)\) has leading coeficient \(a\) and roots \(b\) and \(c\).

OpenStudy (zzr0ck3r):

ya'll makin this way to hard

OpenStudy (zzr0ck3r):

imo

OpenStudy (mathmale):

substitute -4 for x inside the first set of parentheses, and then subst. 6 for x inside the second set of parentheses. Use 5 as the coeff (a) of the highest ordered term (x^2). Multiply out this whole expression and simplify it. Now test x=-4 and x=6. If y our equation is correct, both of these x values should make the equation = 0. Show your work.